OpenStudy (falling_in_katt):

In < RST, RT = 4, ST = 8, and TX bisects < RTS. Which of the following proportions must be true?

OpenStudy (falling_in_katt):

A. \[\frac{ RX }{ ST }=\frac{ TR }{ SX }\]

OpenStudy (falling_in_katt):

B. \[\frac{ TX }{ XR }=\frac{ ST }{ RT }\]

OpenStudy (falling_in_katt):

C. \[\frac{ RX }{ RT }=\frac{ SX }{ ST }\]

OpenStudy (falling_in_katt):

D. \[\frac{ RX }{ RS }=\frac{ XS }{ ST }\]

OpenStudy (lynfran):

c is the correct option